OpenStudy (anonymous):

A particle is projected upwards from level ground at an angle with horizontal such that its range is twice the maximum height reached by it.The angle made by its velocity vector with the horizontal at ahorizontal distance R/4 from the point of projection(R is horizontal range)?

OpenStudy (anonymous):

45 degrees?

OpenStudy (anonymous):

Your answer is in this picture

OpenStudy (anonymous):

s @anusha.p u r corrct but how?

OpenStudy (anonymous):

d=(v^2/g)*sin2theta

OpenStudy (anonymous):

@an can u solve it

OpenStudy (anonymous):

@shahzadjalbani can u help??

OpenStudy (anonymous):

In what

OpenStudy (anonymous):

by telling how to do??

OpenStudy (anonymous):

@lalaly

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

Hey @shakir I'm not really great at this stuff and I'd probably make you fail but @parthkohli is really good with math stuff :)

OpenStudy (anonymous):

Or! @blues she's not on right now but I heard she's a genius when it comes to this type of stuff :3

OpenStudy (anonymous):

Velocity is not given ? why there are two unknown variables in your question :P ?

OpenStudy (anonymous):

it is give R=2Hmax

OpenStudy (anonymous):

and what about R/4

OpenStudy (anonymous):

so R/4 = Hmax/2

OpenStudy (anonymous):

Good, so did you try compare them ?

OpenStudy (anonymous):

s but not getting

OpenStudy (anonymous):

me too ;)

OpenStudy (anonymous):

Firstly let's express what we need to find in formulas

OpenStudy (anonymous):

v0y = v0 sin alpha v0x = v0 cos alpha so v0y/v0x = tan aplha so if you will get values of v0y and v0x at the R/4 point you will easily find the aplha

OpenStudy (anonymous):

v0y is vertical component of initial velocity v0x is horizontal component of initial velocity

OpenStudy (anonymous):

vox is constant so its value doesn't change through all the trajectory and if Beta is initial angle made by velocity vector the its value is vox = vo cos Beta

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Beta is the angle made by Vo and horizon

OpenStudy (anonymous):

So you need to find only voy component at R/4 point

OpenStudy (anonymous):

@ArchiePhysics both angle and initial velocity are not give so how???

OpenStudy (anonymous):

@anusha.p please don't confuse aplha and the other initial angle Aplha is changing angle so its variable And let Beta be the second angle i.e. initial angle so it is some constant value.

OpenStudy (anonymous):

@shakir you don't have a choice you have to use initial angle in your formula or let's see if we can cancel it out

OpenStudy (anonymous):

at R/4 the time interval value is t = t(total)/4 = R/(4vox)

OpenStudy (anonymous):

Ok @anusha.p it is just the matter of convention I said that just to warn that you have to distinct between two angles. One of them is variable the second is constant.

OpenStudy (anonymous):

i gave up!!!!!!!

OpenStudy (anonymous):

There are some equations that you can use for solving this problem. I will make new designations: T = t(total)/2 - time for a particle to rise the top of trajectory or to fall from it. R - range, I think other letters is familiar to you, here R = vox 2 T Hmax = voy T - gT^2/2 voy = gT using previous two equations we get Hmax = gT^2/2 Let's solve that together

OpenStudy (anonymous):

R = vox 2 T?? why

OpenStudy (anonymous):

R= u cos theta *T

OpenStudy (anonymous):

R = vox t(total) <= do you agree with this? T = t(total)/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok I found that v0y/vox = 2 alpha = arctan 2 but it is not correct. Let's see together what I got, may be you will find mistakes

OpenStudy (anonymous):

u-velocity vth wich the object is thrown. ux=initial horizontal velocity=u cos q,where q is the angle of projection. uy=u sin q=initial vertical velocity vx=ux(at any time t since aceleration in horizontal direction is zero) vy=uy-g*t tan t=vy/vx ,t is the angle made by the velocity vector horizontally.

OpenStudy (anonymous):

from previous equations I derived following v0x = R/2T R = 2 Hmax Hmax = gT^2/2 2Hmax = gT^2 R = gT^2 vox = gT^2/2T = gT/2 voy = gT voy/vox=gT/gT/2=2

OpenStudy (anonymous):

I think I got it

OpenStudy (anonymous):

the time is t = T/2

OpenStudy (anonymous):

I got theta =45 degree, but without using R/4 is it wrong ?

OpenStudy (anonymous):

now,given range=2*max height ( u^2sin2q)/g =2*(usinq)^2 /2g. u get tanq=2.(angle of projection found) now,calculating the time required to reach R/4 ucosq=(R/4)/t.(since h.velocity is const) t=R/(4ucosq) gt=usinq/2(by replacing the range formula) Vy=uy-gt=usinq-usinq/2=usinq/2 tant=vy/vx=usinq/(2*ucosq)=tanq/2=1 therefore t=45 degrees!!!!!!!!!

OpenStudy (anonymous):

Vy(T/2) = Voy - g(T/2) = gT/2 - vertical component of velocity at R/4 or at the moment of time T/2 use that instead, so now we get 1 for tan alpha Vy(T/2)/Vox = 1

OpenStudy (anonymous):

The right answer is 45 degrees as arctan 1 = 45 degrees